OpenStudy (anonymous):

Boyle's Law says that the volume of a gas varies inversely with the pressure. When the volume of a certain gas is 10L , the pressure is 112 kPa (kilopascals). What is the volume when the pressure is 160kPa ?

OpenStudy (jhannybean):

Boyle's Law \(\large PV = k \)

OpenStudy (jhannybean):

V varies inversely with P : \(\large V = \frac{k}{P}\)

OpenStudy (anonymous):

ok?

OpenStudy (jhannybean):

V = 10 L , P= 112 kPA : \(\large 10 \ L = \frac{k}{112 \ \ kPA} \) \(\large k = (10 \ L)(112 \ kPA ) = ?\)

OpenStudy (jhannybean):

what do you get?

OpenStudy (anonymous):

Im not sure ? im stuck i have never learned this.

OpenStudy (jhannybean):

what is 10 x 112?

OpenStudy (anonymous):

1120

OpenStudy (jhannybean):

Good. so k = 1120 L kPA..

OpenStudy (anonymous):

so that will be it?

OpenStudy (jhannybean):

Now we're trying to find the volume. \[\large V = \frac{k}{P}\]We're given P = 160 kPA and we found k = 1120 L kPA\[\large V = \frac{1120 \ L \ kPA}{160 \ kPA}\]\[\large V = \frac{1120}{160} \ L\]

OpenStudy (jhannybean):

simplify and find your answer.

OpenStudy (anonymous):

would it be 7?

OpenStudy (jhannybean):

Yep.

OpenStudy (anonymous):

your being a very big help right know thank you very much :)

OpenStudy (jhannybean):

No problem ^^